OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid 135 . This year, he paid 162 . What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

nvr mind

OpenStudy (anonymous):

increment=162-135=27 percentage increment =(27*100)/135